# Measuring Iterative Temporal Reasoning with [Uncaptioned image]Time Puzzles

### v2

@@ Abstract
We introduce [Uncaptioned image]Time Puzzles, [^1] https://github.com/jaaack-wang/Time-Puzzles a constraint-based date inference task for evaluating iterative temporal reasoning. Each puzzle combines factual temporal anchors with (cross-cultural) calendar relations, admits one or multiple valid solution dates, and is algorithmically generated for controlled, dynamic, and continual evaluation. Across 13 diverse LLMs, [Uncaptioned image]Time Puzzles well distinguishes their iterative temporal reasoning capabilities and remains challenging without tools: GPT-5 reaches only 49.3% accuracy and all other models stay below 31%, despite the dataset’s simplicity. Web search consistently yields substantial gains and using code interpreter shows mixed effects, but all models perform much better when constraints are rewritten with explicit dates, revealing a gap in reliable tool use. Overall, [Uncaptioned image]Time Puzzles presents a simple, cost-effective diagnostic for tool-augmented iterative temporal reasoning.
@@
Zhengxiang Wangthanks: Equal contribution. ZW conceived the research idea. ZD designed and implemented the algorithm for dataset generation. ZW executed the LLM experiments. Both authors contributed significantly to drafting and revising the manuscript. Department of Linguistics & IACS Stony Brook University zhengxiang.wang@stonybrook.edu          Zeyu Dong11footnotemark: 1 Department of Applied Math and Statistics Stony Brook University zeyu.dong@stonybrook.edu
@@ 1 Introduction
Time is a fundamental attribute of human societies, and temporal reasoning is a core component of human intelligence. In practice, humans rarely reason about time in isolation or by rote memorization; instead, they routinely rely on external tools such as clocks, calendars, written records, and online resources to support temporal inference and verification. With the growing deployment of tool-using, agentic LLMs across consumer-facing, research, and industrial settings (OpenRouter.ai, 2025), it is both timely and important to evaluate iterative temporal reasoning in tool-augmented contexts.
Despite extensive study, existing temporal reasoning benchmarks remain mismatched with how temporal reasoning is performed in practice. Prior work largely evaluates temporal understanding in static, single-shot settings (Ning et al., 2020; Zhou et al., 2021, 2019; Qin et al., 2021). While recent benchmarks extend evaluation to LLMs and situated contexts (Tan et al., 2023; Chu et al., 2024; Wang and Zhao, 2024; Wei et al., 2025; Chen et al., 2021), they still do not explicitly evaluate tool-augmented, iterative temporal reasoning.
@@
Our main contributions are threefold:
  1. 1.
    We introduce [Uncaptioned image]Time Puzzles, a novel and dynamic benchmark for evaluating iterative temporal reasoning with tools. The benchmark is enabled by a controllable generation algorithm that supports continual evaluation beyond static test sets prone to memorization.
  2. 2.
    We conduct a comprehensive evaluation across 13 diverse LLMs. Although tool access improves accuracy, substantial gaps remain in producing fully consistent solutions, highlighting the need for more reliable tool use. We show that [Uncaptioned image]Time Puzzles serves as a practical diagnostic for progress in tool-augmented iterative temporal reasoning.
  3. 3.
    For reproducibility, we release the code and datasets for generating challenging constraint-based time puzzles.
@@ 2.1 Task Formulation
Let $\mathcal{D}$ denote the set of all possible Gregorian dates, and let $\mathcal{C}$ map a natural language fact $t$ to the set of dates for which $t$ holds, i.e., $\mathcal{C}(t)\subseteq\mathcal{D}$. As illustrated in Figure 1, each puzzle consists of $N$ natural language facts $F=\{t_{1},t_{2},\dots,t_{N}\}$, with an answer set $\mathcal{A}$ that satisfies all constraints jointly:
@@ 2.2 Data Generation
Algorithm
We generate [Uncaptioned image]Time Puzzles instances using the dynamic procedure in Algorithm 1. The algorithm takes as input a set of pre-configured temporal fact templates and a target solution size $M$, and returns a puzzle instance $(F,\mathcal{A})$ with $|\mathcal{A}|=M$.
Generation proceeds as a randomized search-and-find process. We first anchor the puzzle with a real-world event description $t_{kb}$ (often requiring web search) and sample a seed date $d_{\text{seed}}$ that is temporally consistent with $t_{kb}$. We then randomly sample 3-5 more temporal constraints (e.g., calendar relations) until a solver yields an answer set of size $M$. To improve data generation efficiency, the solver prunes the search space using an entropy-based approach (see Appendix A.2 for details). For our datasets detailed below, it takes around 2 seconds to generate each puzzle in a single core CPU.
Temporal Facts
To ensure diversity and realism, we create a diverse taxonomy of temporal facts (Table 2 in Appendix A.2) that supports constraints at three granularities: year, month, and day. Our facts span lesser-known historical events, well-known anchors, and calendar relations (e.g., day-of-week and non-Gregorian calendars), among others.
@@
We construct two conceptually equivalent datasets for controlled evaluation. The default dataset uses implicit constraints (e.g., an event) to target tool-augmented iterative temporal reasoning. The paired dataset rewrites each implicit constraint by spelling out its relevant dates explicitly, isolating pure iterative temporal reasoning. Each dataset contains 600 puzzles, with solution set sizes evenly distributed from 1 to 6.
Input: Target facts $N$, Target solution size $M$, Universal Date Set $\mathcal{D}$
Output: A puzzle instance $(F,\mathcal{A})$
1
2while True do
 // Step 1: Randomly pick a seed date
3 $t_{kb}\leftarrow\text{SelectRandomKnowledgeBasedEvent()}$;
4 $d_{seed}\leftarrow\text{RandomDate}(\mathcal{C}(t_{kb}))$;
5
 // Step 2: Generate Facts
 // See Appendix A.2 for detail
6 $F\leftarrow\textsc{GenerateFact}(t_{kb},d_{seed},N,\mathcal{D})$;
7
 // Step 3: Solve and Validate
 // See Appendix A.2 for detail
8 $\mathcal{A}\leftarrow\textsc{SolvePuzzle}(F,\mathcal{D})$;
9
 // Step 4: Output condition
10 if $|\mathcal{A}|==M$ then
11    return $(F,\mathcal{A})$;
12
13   end if
14
15 end while
Algorithm 1 TimePuzzles Data Generation Process
@@ 3.1 Experimental Setup
We evaluate 13 diverse LLMs from four model families (Table 1) released in 2025. These include four proprietary GPT models (GPT-5 and GPT-4.1 series) OpenAI (2025b) and nine open-weight models: GPT-oss-20B OpenAI et al. (2025) plus instruction/reasoning variants from DeepSeek-V3.2 DeepSeek-AI et al. (2025), Qwen3 Yang et al. (2025), and Ministral-3-8B Mistral AI (2025). See Table 4 in Appendix C for further details.
@@
All models are prompted using zero-shot chain-of-thought (CoT) prompting Kojima et al. (2022) to encourage step-by-step reasoning. The prompt template is provided in Appendix E.
@@
Given our emphasis on precise iterative reasoning, we primarily evaluate performance using exact match (EM) between the predicted and gold sets. We also consider F1 score and Jaccard index, which are highly correlated with EM (Spearman’s $r=0.92$) and differ from EM slightly (see Appendix B.2). We additionally report output token counts as a proxy for reasoning efficiency.
3.2 No Tool Use
Table 1 shows the aggregated tool-less results using both the implicit (default) and explicit constraints.
@@ 3.2 No Tool Use
Despite its low construction cost, Time Puzzles is challenging for tool-less LLMs.
Our dataset is a proof of concept built entirely from easy-to-search factual knowledge that predates each model’s training cutoff or release date. Yet even GPT-5 achieves only 49.3% EM on the dataset, with all other models below 31% EM. Providing explicit dates instead of implicit factual constraints improves performance, but a clear reasoning gap remains: GPT-5 reaches the best EM at 80.7% and it takes 685B parameters for DeepSeek V3.2-R to get only 77.0% EM. Moreover, model performance tends to decrease as the number of solutions grow (Figure 2).
Exact Match (%) Output Tokens
Implicit Explicit Implicit Explicit
GPT-5 49.3 80.7 (+31.4) 4382.4 3416.9 (-965.5)
GPT-5-nano 14.0 57.2 (+43.2) 5271.0 4903.9 (-367.1)
GPT-4.1 30.0 58.8 (+28.8) 671.6 593.0 (-78.6)
GPT-4.1-nano 8.2 40.0 (+31.8) 1116.9 1023.6 (-93.3)
GPT-oss-20B 10.2 43.8 (+33.6) 5601.0 4965.2 (-635.8)
DeepSeek-V3.2 27.0 68.5 (+41.5) 785.4 704.6 (-80.8)
DeepSeek-V3.2-R 33.3 77.0 (+43.7) 8276.5 5821.8 (-2454.7)
Qwen3-4B-Inst 2.7 21.2 (+18.5) 2621.0 2563.5 (-57.5)
Qwen3-4B-Think 5.5 38.7 (+33.2) 9529.9 10037.3 (+507.4)
Qwen3-8B 10.0 42.7 (+32.7) 7603.6 7161.4 (-442.2)
Qwen3-14B 11.3 46.5 (+35.2) 4879.9 4782.3 (-97.6)
Ministral-3-8B-Inst 3.7 12.3 (+8.6) 1801.0 1650.5 (-150.5)
Ministral-3-8B-R 3.8 18.7 (+14.9) 2117.6 1749.4 (-368.2)
Table 1: Exact-match accuracy and average output tokens, averaged over six solution counts, for implicit (default) and explicit constraints in the tool-less setting.
Time Puzzles effectively distinguishes nuanced iterative temporal reasoning capabilities.
In terms of EM, larger models consistently outperform their smaller counterparts (e.g., GPT-X vs. GPT-X-nano; Qwen3 variants), and reasoning models uniformly surpass instruction-tuned ones (e.g., DeepSeek/Ministral pairs). However, scale alone is insufficient: despite its smaller size, Qwen3-4B-Think substantially outperforms Ministral-3-8B-R, and GPT-oss-20B similarly surpasses both while using only 3.6B active parameters.
Success on Time Puzzles hinges on sustained, structured temporal reasoning rather than exhaustive generation.
For example, GPT-5 outperforms all other reasoning models under both implicit and explicit constraints while producing shorter outputs (except the low-performing Ministral-3-8B-R). Similarly, larger Qwen3 models perform better with much fewer output tokens.
3.3 Web Search
Setup
We investigate if web search can close the performance gap between the implicit and explicit constraints. We consider six LLMs: GPT-5, GPT-4.1, GPT-oss-20B, DeepSeek-V3.2-R, Qwen3-14B, and Ministral-3-8B-R. We run GPT-5 and GPT-4.1 with live web search to represent two paradigms: static, one-off retrieval (GPT-4.1) and agentic, iterative retrieval (GPT-5). We then re-use the search results used in the responses of these two models as fixed context for the open-weight LLMs. For cost reasons, we only run web search on a subset of our dataset with solution counts 1, 3, and 5.
Results
Figure 2 shows that web search consistently improves performance across different solution counts, but does not close the performance gap between implicit and explicit constraints. Moreover, the marginal gains decrease as the solution count increases. Notably, GPT-5-style web search yields larger improvements than GPT-4.1-style web search, suggesting that iterative temporal reasoning benefits more from agentic, repeated retrieval than from a single retrieval pass.
Overall, these results indicate that even with access to web search, state-of-the-art LLMs still struggle to reliably retrieve and, more importantly, integrate evidence to solve iterative temporal reasoning problems and fully resolve temporal constraints.
@@ Results
Figure 3: Exact-match accuracy (left) and output tokens (right) for GPT-5 (top) and GPT-4.1 (bottom) under different conditions for single-solution puzzles. Numbers annotated on bars are mean values, beside error bars. +Web enables web search. +CI enables Code Interpreter.
3.4 Code Interpreter
Setup
We evaluate whether enabling Code Interpreter (+CI) improves performance and/or reasoning efficiency under the previously discussed conditions. For cost reasons, we test only GPT-5 and GPT-4.1 on single-solution puzzles, representing static and agentic uses of +CI, respectively.
Results
Figure 3 shows that enabling Code Interpreter (+CI) degrades GPT-5 performance on implicit constraints, both with and without web search, whereas GPT-4.1 benefits from +CI, particularly when combined with web search. In contrast, performance on explicit constraints remains largely unchanged for both models. Moreover, +CI consistently reduces GPT-5 output length, but does not yield similar verbosity reductions for GPT-4.1. Overall, +CI is not a reliable substitute for resolving implicit temporal constraints, though it can improve reasoning efficiency and complement web search for capable non-reasoning models.
@@ 5 Conclusion
We propose [Uncaptioned image]Time Puzzles, a constraint-based date inference task that targets iterative, tool-augmented temporal reasoning. Although puzzles are synthetically generated and easy to verify, they expose persistent failures of current instruction- and reasoning-tuned LLMs to reliably resolve implicit temporal constraints, even with tool access. Overall, [Uncaptioned image]Time Puzzles offers a simple, cost-effective diagnostic for tool-augmented iterative temporal reasoning and can be systematically extended to support more challenging evaluations.

### v3

@@ Abstract
Tool use, such as web search, has become a standard capability even in freely available large language models (LLMs). However, existing benchmarks evaluate temporal reasoning mainly in static, non-tool-using settings, which poorly reflect how LLMs perform temporal reasoning in practice. We introduce [Uncaptioned image] Time Puzzles, a constraint-based date inference task for evaluating iterative temporal reasoning with tools. Each puzzle combines factual temporal anchors with (cross-cultural) calendar relations and may admit one or multiple valid dates. The puzzles are algorithmically generated, enabling controlled and continual evaluation. Across 13 LLMs, even the best model (GPT-5) achieves only 55.3% accuracy without tools, despite using easily searchable facts. While web search improves performance, models perform substantially better when constraints are rewritten with explicit dates, removing the need for factual lookup. These results reveal a gap in reliable tool use for iterative temporal reasoning.
@@
Zhengxiang Wangthanks: Equal contributions. ZW conceived the research idea. ZD designed and implemented the algorithm for dataset generation. ZW designed and executed the LLM experiments. Both authors contributed significantly to drafting and revising the manuscript. ZW revised the second version of this draft. Department of Linguistics & IACS Stony Brook University zhengxiang.wang@stonybrook.edu          Zeyu Dong11footnotemark: 1 Department of Applied Math and Statistics Stony Brook University zeyu.dong@stonybrook.edu
@@ 1 Introduction
Modern LLMs are increasingly deployed with external tools, such as web search and code interpreter, even in free customer-facing settings (OpenRouter.ai, 2025). These capabilities allow models to retrieve factual information and verify intermediate reasoning steps during inference. However, existing temporal reasoning benchmarks largely evaluate models in static, single-shot settings, which do not reflect the iterative, tool-augmented workflows increasingly used in practice.
Prior work on temporal reasoning has focused on static, single-shot settings (Ning et al., 2020; Zhou et al., 2021, 2019; Qin et al., 2021). While recent benchmarks extend evaluation to LLMs and situated contexts (Tan et al., 2023; Chu et al., 2024; Wang and Zhao, 2024; Wei et al., 2025; Chen et al., 2021), they still do not explicitly evaluate tool-augmented, iterative temporal reasoning
@@
We evaluate 13 diverse LLMs on [Uncaptioned image] Time Puzzles under both tool-less and tool-augmented settings. Overall performance remains limited: even the best model (GPT-5) reaches only 55.3% accuracy without tools, despite the use of easily searchable factual anchors. While web search consistently improves results, all models perform better when the same puzzles are rewritten to replace factual anchors with explicit dates. This gap highlights a key weakness in integrating factual lookup with constraint-based, multi-step temporal reasoning.
@@ 2.1 Task Formulation
[Uncaptioned image] Time Puzzles is a constraint-based date inference task: given a set of temporal constraints, the goal is to identify the date(s) that jointly satisfy all constraints. Formally, let $\mathcal{D}$ denote the set of all possible Gregorian dates, and let $\mathcal{C}$ be an oracle function that maps a natural-language temporal constraint $t$ to the set of dates for which $t$ holds, i.e., $\mathcal{C}(t)\subseteq\mathcal{D}$. As illustrated in Figure 1, each puzzle consists of $N$ natural-language constraints $F=\{t_{1},t_{2},\dots,t_{N}\}$. The answer set $\mathcal{A}$ contains all dates that satisfy every constraint:
@@ 2.2 Data Generation
Input Constraints
We create a simple yet diverse taxonomy of temporal constraint types (Table 3 in Appendix A.2). The taxonomy spans both factual anchors (e.g., historical events, presidents in office, zodiac years) and calendar-structural constraints (e.g., months, seasons, weekdays, or day-of-month relations), each restricting candidate dates at the year, month, or day level.
@@
This paper considers only English as a proof of concept, but [Uncaptioned image] Time Puzzles can be readily adapted to other languages. Figure 1 shows an example puzzle with five constraints.
Puzzle Generation
Puzzles are generated algorithmically through randomized constraint sampling. Each puzzle combines one of the 50 real-world historical events we manually curated with additional temporal constraints from the taxonomy (e.g., calendar relations). We retain combinations that yield puzzles with a desired number of valid solution dates. We open-source our code and generated datasets at https://github.com/jaaack-wang/Time-Puzzles.
@@
We construct two conceptually equivalent datasets for controlled evaluation. The default dataset uses implicit constraints (e.g., an event) to target tool-augmented iterative temporal reasoning. The paired dataset rewrites each implicit constraint by spelling out its relevant dates explicitly, thereby removing the need of factual lookup. Each dataset contains 600 puzzles, with solution set sizes evenly distributed from 1 to 6.
We select 100 puzzles per solution size due to cost constraints, as unrestricted web search is prohibitively expensive, given our limited budget. In addition, experiments on a separate set of 600 puzzles yield similar results, supporting the suitability of the current dataset size (see Section 4).
Quality Control
Because all constraints are independently verifiable, the resulting puzzles are both scalable and high quality. We randomly inspected 30 puzzles, where one author independently attempted to solve them using a web browser and verified the correctness of the generated puzzles. This manual inspection also yielded an estimated human baseline accuracy of 80%.
@@ 3.1 Experimental Setup
We evaluate 13 diverse LLMs from four model families (Table 1) released in 2025. These include four proprietary GPT models (GPT-5 and GPT-4.1 series) OpenAI (2025b) and nine open-weight models: GPT-oss-20B OpenAI et al. (2025) plus instruction/reasoning variants from DeepSeek-V3.2 DeepSeek-AI et al. (2025), Qwen3 Yang et al. (2025), and Ministral-3-8B Mistral AI (2025). See Table 4 in Appendix B for further details.
@@
All models are prompted using zero-shot chain-of-thought (CoT) prompting Kojima et al. (2022) to encourage step-by-step reasoning. The prompt template is provided in Appendix C.
Model EM (%) JI (%) F1 (%) # Tks
GPT-5 55.33 58.62 59.52 4126
GPT-5-nano 15.83 19.37 20.31 5224
GPT-4.1 35.17 39.13 40.14 615
GPT-4.1-nano 9.33 13.53 14.84 1098
GPT-oss-20B 10.83 13.51 14.32 5704
DeepSeek-V3.2-R 36.67 39.33 40.08 7862
DeepSeek-V3.2 30.50 32.89 33.51 737
Qwen3-4B-Inst 3.00 3.89 4.14 2408
Qwen3-4B-Think 5.67 6.61 6.89 9701
Qwen3-8B 10.83 13.45 14.26 7341
Qwen3-14B 13.17 16.32 17.36 4733
Ministral-3-8B-Inst 3.50 5.38 6.04 1797
Ministral-3-8B-R 4.27 4.81 5.04 2058
Table 1: Exact-match accuracy (EM), Jaccard index (JI), F1 score (F1), and output tokens (Tks), averaged over six solution counts, for the 600 generated puzzles with implicit (default) constraints in the tool-less setting.
@@
Let $Y_{i}$ denote the gold solution set and $\hat{Y}_{i}$ the predicted set for instance $i$. We compute exact match (EM), F1 (F1), and Jaccard index (JI):
$EM_{i}=\mathbb{I}[Y_{i}=\hat{Y}_{i}].$ (2)
$P_{i}=\tfrac{|Y_{i}\cap\hat{Y}_{i}|}{|\hat{Y}_{i}|},\;R_{i}=\tfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}|},\;F1_{i}=\tfrac{2P_{i}R_{i}}{P_{i}+R_{i}}.$ (3)
$JI_{i}=\begin{cases}1&Y_{i}=\hat{Y}_{i}=\emptyset\\ \tfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}\cup\hat{Y}_{i}|}&\text{otherwise}.\end{cases}$ (4)
Given our emphasis on precise iterative reasoning, we treat EM as the primary metric.
Tools
We consider the following tool-augmented settings: web search, Code Interpreter (+CI), and their combinations. Due to cost constraints, these experiments are conducted on puzzles with solution counts 1, 3, and 5 and with selected models.
More specifically, for web search, we test six best-performing LLMs across different model families (GPT-5, GPT-4.1, GPT-oss-20B, DeepSeek-V3.2-R, Qwen3-14B, and Ministral-3-8B-R), whereas for +CI, we only evaluate GPT-5 and GPT-4.1 both with and without web search. We use off-the-shelf tools via OpenAI Response API and we include GPT-5 and GPT-4.1 across all tool-augmented settings to represent tool-using capablities by a reasoning model (GPT-5) as well as an instruction model (GPT-4.1). Note that GPT-5 and GPT-4.1 are run with live web search and their retrieved results are reused as fixed context for the open-weight models.
3.2 Results
@@
Despite its simplicity, [Uncaptioned image] Time Puzzles is challenging for tool-less LLMs.
Our dataset is a proof of concept built from easy-to-search facts that predates each model’s training cutoff or release date. Yet even GPT-5 achieves only 55.3% EM on the dataset, with all other models below 37% EM (Table 1). JI and F1 yield slightly higher scores, but the overall performance patterns remain consistent, with a Spearman correlation of $r=0.93$. Moreover, the three metrics agree in 88.6% of cases, indicating that LLMs typically either solve a puzzle completely or fail to solve it at all.
Tool use, particularly web search, can greatly improve model performance, but a clear reasoning gap remain.
Figure 2 shows that web search consistently improves performance across different solution counts, but models nearly always perform the best when the implicit constraints in [Uncaptioned image] Time Puzzles are rewritten with explicit dates to remove the need for factual lookup. Similarly, Figure 3 shows that enabling Code Interpreter (+CI) does not close the gap either: +CI degrades GPT-5 performance on implicit constraints, both with and without web search. While GPT-4.1 benefits from +CI when combined with web search, the performance is below that of explicit constraints.
@@
Figure 3: Average EM accuracy (%) for GPT-5 (left) and GPT-4.1 (right) under different conditions across three solution counts (1, 3, and 5). Numbers annotated on bars are mean values, beside error bars. +Web enables web search. +CI enables Code Interpreter.
4 Analysis
Section 3 shows that [Uncaptioned image] Time Puzzles provides a challenging and cost-effective benchmark for iterative temporal reasoning. This section further demonstrates that as a benchmark, [Uncaptioned image] Time Puzzles is both discriminative and reliable.
[Uncaptioned image] Time Puzzles effectively distinguishes nuanced iterative temporal reasoning capabilities.
Table 1 clearly shows that larger models consistently outperform their smaller counterparts (e.g., GPT-X vs. GPT-X-nano; Qwen3 variants), and reasoning models uniformly surpass instruction-tuned ones (e.g., DeepSeek/Ministral pairs). However, scale alone is insufficient: despite its smaller size, Qwen3-4B-Think substantially outperforms Ministral-3-8B-R, and GPT-oss-20B similarly surpasses both while using only 3.6B active parameters. Moreover, for reasoning models, higher performance typically uses less output tokens, suggesting that success on [Uncaptioned image] Time Puzzles hinges on sustained, structured reasoning rather than exhaustive generation (e.g., Qwen3 models in Table 1).
Our findings are robust to prompt variations and puzzle sets.
Table 2 shows that inserting a detailed step-by-step reasoning instruction into the default prompt (see Appendix C) decreases performance, particularly for larger models. Furthermore, repeating the experiments from Section 3 on 600 newly generated puzzles yields negligibly different exact-match (EM) scores, validating the suitability of our evaluation dataset size.
Type Setting G4.1-Nano G5-Nano G4.1 G5
Implicit Default 9.3 15.8 35.2 55.3
+Reason Inst 8.8 14.2 25.3 47.3
New Runs 11.0 16.8 29.8 50.1
Explicit Default 41.8 60.2 63.0 84.7
+Reason Inst 35.5 55.5 53.5 77.7
New Runs 42.7 60.2 63.7 80.5
Table 2: EM accuracy (%) averaged over six solution counts under different settings for both explicit and implicit constraint across the four GPT models. Default setting is based on Section 3. “+Reason Inst” enhances the default setting with detailed step-by-step reasoning instruction. “New Runs” reruns the default setting with completely new puzzles generated from scratch.
@@ 6 Conclusion
We propose [Uncaptioned image] Time Puzzles, a constraint-based date inference task that targets iterative, tool-augmented temporal reasoning. Although puzzles are synthetically generated and easy to verify, they are challenging for tool-less LLMs. They also expose persistent failures of current instruction- and reasoning-tuned LLMs to reliably resolve implicit temporal constraints, even with tool access. Overall, [Uncaptioned image] Time Puzzles offers a simple, cost-effective, and discriminative diagnostic for tool-augmented iterative temporal reasoning and can be systematically extended to support more challenging evaluations or additional languages (e.g., via translation).
@@
Language Diversity
Our work currently only considers English language, due to budget constaints. Future studies can extend our study by adapting our evaluation framework to other languages. This can be easily done by, for example, converting the puzzle template into other languages.